OpenStudy (anonymous):

Circle S has an equation of (x – 16)2 + (y – 9)2 = 4. What is the center and radius of circle S?

OpenStudy (anonymous):

General equation of the circle is given by: \[\large (x-h)^2 + (y-k)^2 = r^2\] Here (h,k) is center and r is radius.. Can you compare for the values of h k and r??

OpenStudy (anonymous):

You can write your equation as: \[\large (x-16)^2 + (y-9)^2 = 2^2\] Now compare it with general equation and tell me the values of h, k and r...

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

9^2 = 81 and is the radius?

OpenStudy (anonymous):

Im not comprehending.

OpenStudy (anonymous):

Right hand side 4..

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

I see it!

OpenStudy (anonymous):

See tell me what I have written and what your equation is can you see they are similar???

OpenStudy (anonymous):

They are both negative. I got Center: (-16, -9); Radius: 4

OpenStudy (anonymous):

I show you more clearly: \[\large (x-h)^2 + (y-k)^2 = r^2\] \[\large (x-16)^2 + (y-9)^2 = 2^2\] They are similar or not??

OpenStudy (anonymous):

They are you plug it in.

OpenStudy (anonymous):

no they are not negative.. if you closely look then h is 16 and not -16.. look it once more.. because there is minus sign already in front of it..

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):

So its positive? Center: (16, 9); Radius: 4

OpenStudy (anonymous):

Radius in not 4.. Check it once again...

OpenStudy (anonymous):

Is it 2?

OpenStudy (anonymous):